OpenStudy (anonymous):

Which answer shows this equation written in standard form? 4 – 3(x – y) = –8x + 2 A. –11x + 3y = –2 B. 5x – 3y = 2 C. 5x + 3y = –2 D. 5x + y = 2

OpenStudy (anonymous):

@KamiBug plz help

OpenStudy (danjs):

standard linear form

OpenStudy (danjs):

Ax + By = C

OpenStudy (danjs):

where A,B, and C are constants, numbers

OpenStudy (danjs):

so 4-3(x-y) = -8x + 2

OpenStudy (danjs):

you need to put that into the Ax + By = C form.

OpenStudy (danjs):

multiply out the -3(x-y) first

OpenStudy (anonymous):

umm....I did it and I think it is C. Am I right?

OpenStudy (danjs):

i dont know the answer , all i did so far was 4 - 3x + 3y = -8x + 2

OpenStudy (danjs):

Now add 8x to both sides

OpenStudy (danjs):

5x +3y +4 = 2

OpenStudy (danjs):

now take away 4 from both ends

OpenStudy (danjs):

5x + 3y = -2

OpenStudy (danjs):

Which part of it were you unsure about ?

OpenStudy (anonymous):

sorry I just got bit confused thanks though

OpenStudy (anonymous):

The standard for for a line: Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.